40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 28 02:43:49 2013, yuta, Update, LSC, some qualitative evidence of PRMI sideband lock PRMIsideband.png
    Reply  Thu Feb 28 03:25:27 2013, yuta, Update, LSC, PR gain ~ 25 from PRMI carrier lock PRMIcarrier.png
       Reply  Thu Feb 28 03:41:31 2013, Koji, Update, LSC, PR gain ~ 25 from PRMI carrier lock 
       Reply  Sat Mar 2 05:53:15 2013, yuta, Update, LSC, stable lock of PRMI LSCMICHOLTF_PRMI.pngLSCPRCLOLTF_PRMI.pngPRMtoAS55REFL55.png
          Reply  Sat Mar 2 14:52:02 2013, rana, Update, LSC, stable lock of PRMI 
             Reply  Sun Mar 3 22:16:46 2013, Jenne, Update, LSC, LSC whitening triggering started 
                Reply  Mon Mar 4 09:55:33 2013, rana, Update, LSC, LSC whitening triggering started 
                Reply  Tue Mar 5 18:36:27 2013, Jenne, Update, LSC, LSC whitening triggering started 
Message ID: 8212     Entry time: Sat Mar 2 05:53:15 2013     In reply to: 8197     Reply to this: 8213
Author: yuta 
Type: Update 
Category: LSC 
Subject: stable lock of PRMI 

I tuned alignment, gains and filters to achieve stable lock of PRMI.
It now locks quite stably with UGF of ~100 Hz. Measured power recycling gain at maximum is ~ 25.

Locking details:
  == PRMI carrier ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = -12 deg,  MICH_GAIN = -1, feedback to ITMX(-1),ITMY(+1)
  PRCL: REFL55_I_ERR, REFL55_PHASE_R = 70 deg, PRCL_GAIN = 0.3, feedback to PRM

  MICH servo is always on. PRCL loop turns on by trigger using POP DC. Boost filters and resonant gains turn on by triggers using POP DC.
  Gain normalization was not used.


Openloop transferfunctions:

  MICH: UGF ~90 Hz, phase margin ~40 deg
  PRCL: UGF ~100 Hz, phase margin ~50 deg (cf. Fitted gain was same as half-PRC: elog #8053)
LSCMICHOLTF_PRMI.png    LSCPRCLOLTF_PRMI.png



Power recycling gain:

  POP DC when unlocked is 6, when locked is 2200-2500, and when dark is 0. So, power recycling gain is ~ 22 to 25. Value without any loss in PRMI is 45 (elog #6947). Alignment was pretty critical to achieve this recycling gain and stable lock.
  There was oscillation at 630 Hz when locked, which is similar to the one we saw in POX11 (elog #8203).


Youtube:





AS(top left), POP(top right), REFL(bottom left), and ETMYT(bottom right). ETMY was mis-aligned, but you can see the beam at ETMYT after PRMI was carrier locked.



MICH/PRCL coupling:

  I measured "sensing matrix" of PRMI by tickling PRM/ITMs/BS at 8.5 Hz and measuring 8.5 Hz peak height of AS55_Q, REFL55_I spectra during PRMI lock (attached is an example measurement of PRM). Below table is the result. AS55_Q has ~5% of sensitivity to PRCL compared with MICH. Also, BS introduces REFL55_I signal considerably. And also, there seems to be an imbalance in actuation efficiency between ITMX and ITMY.

actuation AS55_Q_ERR   REFL55_I_ERR
ITMX      +11.4        +0.80
ITMY      +33.0        +1.06
BS        +50.8        +1.90
PRM       - 0.7        +1.05



AS clipping:
  AS was clipped inside the vaccuum the other day(elog #8198). So, I tried to avoid AS clipping by aligning BS this morning. But it turned out that avoiding AS clipping by BS makes ITMX beam centering worse. Maybe we should center the beam on Yarm first next week.


Next:
 - calculate expected PRMI recycling gain with loss, PR2 filpped
 - beam centering on the arms
    - IPANG, IPPOS, Y green, X green
    - PRMI g-factor measurement

Attachment 3: PRMtoAS55REFL55.png  55 kB  | Hide | Hide all
PRMtoAS55REFL55.png
ELOG V3.1.3-